# Critical Metallicity of Cool Supergiant Formation. II. Physical Origin

### v1

@@ 5.2 Homology Relation between Metallicity and Radius
The effects of opacity and nuclear energy generation rates on stellar radius can be understood through homology relations (see, e.g., Maeder, 2009; Kippenhahn et al., 2012). The central idea of the homology method is to compare two stars with similar internal structures. At a given relative mass coordinate, defined as $\xi=m/M$, where $M$ is the total stellar mass and $m$ is the local mass coordinate, the fundamental physical quantities in the two stars can be transformed with simple scaling constants. In the analysis below, we simply use the global stellar parameters such as $M$ or the average values within the entire star to express the homology relations to avoid complexity.
@@
When $Z\gg 10^{-3}$, bound-free absorption dominates and thus $\kappa_{0}\sim Z$; when $Z\ll 10^{-3}$, free-free absorption dominates, and thus the opacity becomes largely independent of $Z$. Near the critical metallicity $Z\sim 0.001$ for RSG formation, both contributions are comparable. The coefficient $\epsilon_{0}$ for the burning of the CNO cycle is proportional to the mass fraction of carbon, nitrogen, and oxygen, and we approximate it as $\epsilon_{0}\approx Z$. For a fully ionized ideal gas, the mean molecular weight is approximated by:
$\mu\simeq\frac{4}{3+5X-Z}$ (13)
Assuming a fixed He mass fraction $Y=0.278$ in our models, the H mass fraction becomes $X=0.722-Z$, and thus:
$\mu^{-1}\simeq\frac{6.61-6Z}{4}$ (14)
. Substituting these $Z$ dependencies into Equation (11), we obtain:
@@
We define the coefficient dependent on $Z$ as $C_{Z}\equiv[Z(1+10^{3}Z)]^{0.051}(1.1-Z)^{-0.49}$, with contributions from nuclear burning $C_{\epsilon}\equiv Z^{0.051}$, opacity $C_{\kappa}\equiv(1+10^{3}Z)^{0.051}$, and mean molecular weight $C_{\mu}=(1.1-Z)^{-0.49}$. The values of these coefficients as functions of $Z$ are plotted in Figure 12. At $Z=2\times 10^{-2}$, the value of $C_{Z}$ is $\sim 1.7$ times greater than at $Z=10^{-5}$. Although this difference is somewhat smaller than that seen in the TAMS radii between these two metallicities in Figure 8 ($\sim 2-3$ times difference between $Z=2\times 10^{-2}$ and $Z=10^{-5}$), the expression nonetheless captures the general trend of increasing the radius with metallicity.
Among the three components, the contribution from mean molecular weight is relatively mild, as $(1.1-Z)^{-0.49}$ remains close to unity and varies only slightly across the metallicity range considered. For $Z\lesssim 10^{-3}$, the $C_{\kappa}$ also stays near unity, making $C_{\epsilon}$ the dominant contributor to $C_{Z}$. At higher metallicities ($Z\gtrsim 10^{-3}$), the contribution from opacity becomes more significant, while $C_{\epsilon}$ continues to play a role. This trend is consistent with the results from Grid (b) in Figure 1 and Grid (c) in Figure 2, which show that stellar radii become sensitive to opacity only when $\zeta_{\rm opacity}\gtrsim 10^{-3}$. Here, $\zeta_{\rm opacity}$ serves as a pseudo-metallicity parameter used for interpolating opacities in the opacity tables, effectively mimicking the metallicity dependence of opacity. Overall, the increase in stellar radius with metallicity during the main sequence is primarily due to enhanced opacity and more efficient energy generation through the CNO cycle.
We now examine the mass dependence of the stellar radius at fixed metallicity. From the homology relation, we obtain $R\sim M^{0.69}$ for the CNO cycle. This exponent can be compared to the threshold radius for RSG formation across different stellar masses, as shown in Figure 8, which follows a rough power-law scaling of $R\sim M^{0.75}$. Although this exponent 0.75 is only a coarse fit, it matches closely to the exponent 0.69 of the constant-metallicity relation. This similarity indicates that the the mass dependence of the threshold radius for RSG formation is closely aligned with mass dependence of stellar radius at constant metallicity. Consequently, the existence of a nearly constant critical metallicity at $Z\sim 0.001$ across a range of masses can be well justified.
@@
6 Classification of Supergiant Evolution Pathways
To understand the origin of the threshold TAMS radius that determines RSG formation, we build on the evolutionary pathways proposed in our previous work Ou & Chen (2024). In that study, we identified two pathways through which stars evolve into the RG/RSG phase, both governed by a refined form of the “mirror principle,” wherein the stellar envelope expands or contracts in the opposite direction of the shell-envelope boundary. Initially, as the shell-envelope boundary recedes inward during core contraction, the envelope expands accordingly. In some models, the envelope expands into the RG/RSG regime during this phase—this represents the first evolutionary pathway to RG/RSG formation.
After core contraction ends, the nuclear energy generation rate ($L_{\rm nuc}$), which includes contributions from shell H burning and core He burning, dominates the energy production of the stellar interior. Although $L_{\rm nuc}$ dominates the energy source during this phase, it is already in decline. As $L_{\rm nuc}$ decreases, the shell-envelope boundary continues to move inward, causing the envelope to expand further. If $L_{\rm nuc}$ subsequently increases after reaching a minimum, the envelope reverses and contracts, and the star remains a BSG. If, on the other hand, $L_{\rm nuc}$ continues to decline monotonically, the envelope continues to expand into the RSG regime, this represents the second pathway to the evolution of RG / RSG.
@@ 6 Classification of Supergiant Evolution Pathways
Based on these evolutionary pathways, we classify each massive star model in Grid (a). In addition to the above scenarios, we identify an additional evolutionary type that occurs in stars with initial masses $\gtrsim 30\,M_{\odot}$ and metallicities $Z\lesssim 0.001$. In these cases, the C core begins to form while the He core is still contracting. The envelope may or may not expand further after the onset of core C burning, and this complex evolution in the final stages lies beyond the scope of this study. As our classification focuses on the core He-burning stage, we categorize these stars as BSGs on the basis of their status before the formation of the C core.
@@
Combining all cases, we classify the core He-burning evolution of massive stars into four distinct types:
@@
    RSG-L: The luminosity of nuclear burning ($L_{\rm nuc}$) undergoes a sustained decline after the core contraction, driving a continued expansion of the envelope in the RSG regime.
@@
    BSG-L: The luminosity of nuclear burning ($L_{\rm nuc}$) increases shortly after core contraction, causing the envelope to contract, and the star remains a BSG.
@@
If $R_{P\rm max}$ and $R_{\rm max}$ differ by less than 1%, this indicates that the maximum radius is reached while the central pressure is still increasing, indicating that the core is contracting at that time. For stars that satisfy this condition and have an effective temperature $\log(T_{\rm eff}/{\rm K})<3.8$ at $R_{\rm max}$, we classify them as RSG-C.

### v2

@@ V.2 Homology Relation between Metallicity and Radius
The effects of opacity and nuclear energy generation rates on stellar radius can be understood through homology relations (see, e.g., Maeder, 2009; Kippenhahn et al., 2012). The central idea of the homology method is to compare two stars with similar internal structures. At a given relative mass coordinate, defined as $\xi=m/M$, where $M$ is the total stellar mass and $m$ is the enclosure mass, the fundamental physical quantities in the two stars can be transformed with simple scaling constants. In the analysis below, we simply use the global stellar parameters such as $M$ or the average values within the entire star to express the homology relations to avoid complexity.
@@
When $Z\gg 10^{-3}$, bound-free absorption dominates and thus $\kappa_{0}\sim Z$; when $Z\ll 10^{-3}$, free-free absorption dominates, and thus the opacity becomes largely independent of $Z$. Near the critical metallicity $Z\sim 0.001$ for RSG formation shown in the models of Paper I, both contributions are comparable. The coefficient $\epsilon_{0}$ for the burning of the CNO cycle is proportional to the mass fraction of carbon, nitrogen, and oxygen, and we approximate it as $\epsilon_{0}\approx Z$. For a fully ionized ideal gas, the mean molecular weight is approximated by:
$\mu\simeq\frac{4}{3+5X-Z}.$ (13)
Assuming a fixed He mass fraction $Y=0.278$ in our models, the H mass fraction becomes X=0.722ZX=0.722-Z, and thus:
$\mu^{-1}\simeq\frac{6.61-6Z}{4}.$ (14)
Substituting these $Z$ dependencies into Equation (11), we obtain:
@@
We define the coefficient dependent on $Z$ as $C_{Z}\equiv[Z(1+10^{3}Z)]^{0.051}(1.1-Z)^{-0.49}$, with contributions from nuclear burning $C_{\epsilon}\equiv Z^{0.051}$, opacity $C_{\kappa}\equiv(1+10^{3}Z)^{0.051}$, and mean molecular weight $C_{\mu}=(1.1-Z)^{-0.49}$. The values of these coefficients as functions of $Z$ are plotted in Figure 10.
Among the three components, the contribution from mean molecular weight is relatively mild, as $(1.1-Z)^{-0.49}$ remains close to unity and varies only slightly across the metallicity range considered. For $Z\lesssim 10^{-3}$, the $C_{\kappa}$ also stays near unity, making $C_{\epsilon}$ the dominant contributor to $C_{Z}$. At higher metallicities ($Z\gtrsim 10^{-3}$), the contribution from opacity becomes more significant, while $C_{\epsilon}$ continues to play a role. Overall, the increase in stellar radius with metallicity during the main sequence is primarily due to enhanced opacity and more efficient energy generation through the CNO cycle.
At fixed metallicity, Equation (15) gives $R\propto M^{0.69}$. We compare this scaling with the threshold TAMS radius for RSG formation across different stellar masses, shown in Figure 6, which follows an approximate relation $R\propto M^{0.75}$. Although the latter exponent (0.75) is only a coarse fit, it is close to the value (0.69) expected at constant metallicity. This similarity indicates that the mass dependence of the threshold TAMS radius is closely aligned with the mass dependence of stellar radius at fixed metallicity. As a result, the near constancy of the critical metallicity at $Z\sim 0.001$ over a wide mass range shown in the models of Paper I can be naturally explained.
@@
VI Impact of TAMS Radius on Supergiant Expansion
The remaining key question is how $R_{\rm TAMS}$ regulates the subsequent supergiant expansion, and establishes a threshold for RSG formation. To address this, we first use the $25\,M_{\odot}$ models in Grids (c) and (a) to show the key role of $R_{\rm TAMS}$, and then classify each model in the full Grid (a) to examine the evolution of stars with different masses.
VI.1 Role of TAMS Radius for $25\,M_{\odot}$ Stars
Variations in $\eta_{\rm CNO}$ and $\zeta_{\kappa}$ in Grid (c) produce different values of $R_{\rm TAMS}$. To examine how $R_{\rm TAMS}$ influences later evolution, we select several example models from this grid for comparison. Figure 11 shows their evolutionary tracks in the $(R_{*},P_{\rm c})$ plane, where $R_{*}$ traces the envelope extent and $P_{\rm c}$ reflects the core conditions. All models follow a similar pattern in core evolution: $P_{\rm c}$ rises rapidly, enters a brief plateau, and then increases again. However, their tracks are systematically shifted along the $R_{*}$ axis, and this horizontal offset correlates strongly with $R_{\rm TAMS}$ rather than with the individual input parameters $\eta_{\rm CNO}$ or $\zeta_{\kappa}$. Thus, although all models undergo comparable core-evolution stages, those with larger $R_{\rm TAMS}$ reach each stage at correspondingly larger envelope radii.
To explain how this systematic offset influences the fate of supergiants, we present the evolution of $L_{\rm nuc}$ for the same set of $25\,M_{\odot}$ models in Figure 12. For these models, the bifurcation into RSG or BSG outcomes follows the criterion of Ou & Chen (2024): after core contraction ends, if $L_{\rm nuc}$ turns upward while the star is still in the BSG regime, envelope expansion halts and the model remains a BSG; conversely, if $L_{\rm nuc}$ continues to decline, the envelope expands into the RSG phase. Decomposing $L_{\rm nuc}$ into contributions from H burning ($L_{\rm H}$) and He burning ($L_{\rm He}$), we find that the key driver of the $L_{\rm nuc}$ upturn in BSG cases is the second rise of $L_{\rm He}$. This second rise occurs in all models. However, if it takes place only after the star has already expanded into the stable RSG regime, it has little influence on the envelope. In contrast, if the same increase occurs before the star enters the RSG phase, it can halt further expansion and cause the star to remain a BSG.
The crucial role of $R_{\rm TAMS}$ lies in setting the envelope radius at the evolutionary stage corresponding to the second rise of $L_{\rm He}$. A star with a larger $R_{\rm TAMS}$ maintains a systematically larger envelope radius at all subsequent core-evolution stages, including at the time of the second rise of $L_{\rm He}$. The timing of this second rise relative to the envelope expansion then determines whether the star ultimately ends its evolution as a BSG or an RSG.
We now examine the $25\,M_{\odot}$ models in Grid (a) across different metallicities by plotting the evolution of $L_{\rm nuc}$, $L_{\rm H}$, and $L_{\rm He}$ as functions of $R_{*}$, as shown in Figure 13. Lower-metallicity stars have smaller $R_{\rm TAMS}$ and therefore encounter the second rise of $L_{\rm He}$ at an earlier evolutionary stage, which prevents them from expanding into the RSG regime. In contrast, higher-metallicity stars enter the RSG phase before the second rise of $L_{\rm He}$ occurs. This behavior is consistent with the mechanism identified in Grid (c) (Figure 12).
Figure 11: Evolution of central pressure ($P_{\rm c}$) as a function of stellar radius ($R_{*}$) during the core He-burning stage for $25\,M_{\odot}$, $Z=0.001$ models with varying $\eta_{\rm CNO}$ and $\zeta_{\kappa}$. The legend is ordered by the TAMS radius ($R_{\rm TAMS}$) of each model. Notably, the horizontal shifts in $R_{*}$ among the evolutionary tracks align with $R_{\rm TAMS}$, rather than with the individual input parameters $\eta_{\rm CNO}$ or $\zeta_{\kappa}$.
Figure 12: Evolution of the nuclear luminosity $L_{\rm nuc}$ and its contributions from H burning ($L_{\rm H}$) and He burning ($L_{\rm He}$) as functions of stellar radius ($R_{*}$) for $25\,M_{\odot}$ models shown in Figure 11. Red curves denote models with larger $R_{\rm TAMS}$ that evolve into the RSG phase, while blue curves denote models with smaller $R_{\rm TAMS}$ that remain in the BSG phase. To reduce visual clutter, the tracks are truncated at the point where $L_{\rm He}$ reaches its maximum. Arrows indicate the direction of evolution.
Figure 13: Evolution of the nuclear luminosity $L_{\rm nuc}$ and its contributions from H burning ($L_{\rm H}$) and He burning ($L_{\rm He}$) as functions of stellar radius ($R_{*}$) for $25\,M_{\odot}$ models with different metallicities from Grid (a). Red curves denote higher-metallicity models (and thus larger $R_{\rm TAMS}$) that evolve into the RSG phase, while blue curves denote lower-metallicity models (with smaller $R_{\rm TAMS}$) that remain in the BSG phase. To reduce visual clutter, the tracks are truncated at the point where $L_{\rm He}$ reaches its maximum. Arrows indicate the direction of evolution.
VI.2 Classification of Supergiant Evolution Pathways
To develop a general explanation for the critical metallicity across different stellar masses, we examine the supergiant evolutionary pathways of all the models in Grid (a). We classify the models following Ou & Chen (2024), in which we identified two routes for stars to enter the RG/RSG phase, both governed by the refined “mirror principle.” The first is direct envelope expansion during core contraction, referred to here as the RSG-C case. The second is continued expansion after core contraction ends, driven by the decline of the nuclear luminosity $L_{\rm nuc}$—the combined contribution of shell H burning and core He burning—which we call the RSG-L case. In contrast, if $L_{\rm nuc}$ rises again before turning into an RG/RSG, the envelope re-contracts and the star remains a BSG; we denote this as the BSG-L case.
@@ VI.2 Classification of Supergiant Evolution Pathways
In addition to the above cases, there exists an additional type for stars with initial masses $\gtrsim 30\,M_{\odot}$ and metallicities $Z\lesssim 0.001$ that the C core begins to form while the He core is still contracting. The envelope may or may not expand further after the onset of core C burning, and this complex evolution in the final stages lies beyond the scope of this study. We categorize these stars as BSG-E on the basis of their status before the formation of the C core.
@@
We classify the core He-burning evolution of massive stars into four distinct types:
@@
    RSG-L: $L_{\rm nuc}$ undergoes a sustained decline after the core contraction, driving a continued expansion of the envelope in the RSG regime.
@@
    BSG-L: $L_{\rm nuc}$ increases shortly after core contraction, causing the envelope to contract, and the star remains a BSG.
@@
The detailed classification criterion for each type is described in Appendix A.
VI.3 TAMS Radius and the Critical Metallicity
The classification results are shown in Figure 14 and compared with the maximum supergiant radii. The critical metallicity for cool-supergiant formation at $Z\sim 0.001$ shown in Grid (a) does not correspond to a single evolutionary boundary. Instead, it aligns with different pathway transitions in different mass ranges: it marks the boundary between the RSG-L and BSG-L cases for $M_{i}\lesssim 30\,M_{\odot}$, but the boundary between the RSG-C and BSG-E cases for $M_{i}\gtrsim 30\,M_{\odot}$.
[…]
The transition between RSG-C and BSG-E cases for stars with $M_{i}\gtrsim 30\,M_{\odot}$ can likewise be interpreted in terms of $R_{\rm TAMS}$. For models with smaller $R_{\rm TAMS}$, advanced core-evolution stages occur while the envelope is still relatively compact. The BSG-E cases are those in which the star proceeds into core C burning while the He core is still contracting and the envelope is still expanding in the BSG regime.
[…]
VI.4 Correspondence between core and envelope evolution
The bifurcation between RSG and BSG outcomes arises from how the evolutionary stages of the core and the envelope correspond to one another. As summarized in the flow chart in Figure 15, the post–main-sequence evolution of the core and the envelope is not synchronized; each follows its own sequence of structural transitions. After the TAMS, the core begins to contract. When core contraction ends, $L_{\rm nuc}$ first decreases and then increases again due to the second rise of $L_{\rm He}$. A more advanced core stage is reached once the C core forms. Meanwhile, the envelope initially expands into the BSG regime and may later continue expanding into the RSG regime.
The importance of $R_{\rm TAMS}$ is that it is the stellar radius at the onset of post–main-sequence evolution, serving as the initial condition for subsequent envelope expansion, and thus leading to different evolutionary pathways. A larger $R_{\rm TAMS}$ allows the star to reach large radii earlier during its expansion, so that it is already in the stable RSG state when the second rise of $L_{\rm He}$ occurs (the RSG-C and RSG-L cases). Conversely, a smaller $R_{\rm TAMS}$ keeps the star compact until a later core-evolution stage, causing the second rise of $L_{\rm He}$ to occur while the star is still in the BSG regime (BSG-L cases), or—at higher masses—to proceed quickly into core C burning while still compact (BSG-E cases).
Figure 15: Flow chart illustrating the evolutionary stages of the stellar core and envelope. The stellar radius at the TAMS is the key factor that determines the evolutionary pathway and the resulting supergiant type.
VII Discussions and Conclusions
This study investigates the physical origin of the critical metallicity for RSG formation through a series of numerical experiments. We show that metallicity determines the supergiant outcome primarily through its influence on the stellar radius. In particular, $R_{\rm TAMS}$—strongly shaped by metallicity—defines a threshold for RSG formation, with the threshold value depending on stellar mass.
Metallicity significantly influences stellar evolution through its effects on opacity, nuclear reaction rates, and the mean molecular weight. Using stellar models together with dimensional analysis, we show that higher-metallicity stars tend to reach larger $R_{\rm TAMS}$ due to enhanced opacity and increased CNO-cycle energy generation. The resulting $R_{\rm TAMS}$ then serves as the initial condition for envelope expansion during post–main-sequence evolution.
The mechanism linking $R_{\rm TAMS}$ to the supergiant outcome lies in the correspondence between the envelope radius and the core evolutionary stage. A larger $R_{\rm TAMS}$ means that post–main-sequence expansion begins from a larger initial radius, allowing the envelope to reach larger sizes at subsequent core-evolution phases. For stars with $M_{\rm i}\sim 11$$30\,M_{\odot}$, $R_{\rm TAMS}$ determines whether the envelope has already expanded to RSG dimensions by the time of the second rise of $L_{\rm He}$, after which further expansion is suppressed. For more massive stars with $M_{\rm i}\sim 30$$60\,M_{\odot}$, a smaller $R_{\rm TAMS}$ can lead the star to enter core carbon burning while the envelope remains compact in the BSG phase. The evolution in this higher-mass regime is more complex and requires further investigation.
In addition to metallicity, other stellar properties—such as rotation and mass loss—have been suggested to influence whether a star evolves into a red or blue supergiant (e.g., He et al., 2025). Our results imply that any parameter capable of modifying $R_{\rm TAMS}$ can affect the supergiant outcome. Consequently, one may expect not only a critical metallicity but also threshold values in other physical parameters for RSG formation, which will be explored in future studies.
It is worth noting that the supergiant outcome is sensitive to mixing prescriptions in stellar models (e.g., Langer & Maeder, 1995; Sibony et al., 2023). In particular, recent studies by Schootemeijer et al. (2019) showed that enhanced overshooting and reduced semi-convection tend to favor evolution toward RSGs rather than BSGs. To assess the impact of these mixing parameters, we present additional grids of models in Appendix B, in which the overshooting parameter $f_{\rm ov}$ and the semi-convection efficiency $\alpha_{\rm sc}$ are modified relative to Grid (a). We find that while $f_{\rm ov}$ and $\alpha_{\rm sc}$ influence the detailed supergiant population, the critical metallicity at $Z\sim\text{a few}\times 10^{-3}$ is largely unchanged. Furthermore, for certain values of the convective overshooting parameter, stellar evolutionary tracks may undergo a “blue loop" after entering the RSG phase (e.g., Wagle et al., 2019, 2020). This behavior is beyond the scope of the present study and warrants further investigation to obtain a more complete picture of BSG/RSG evolution.
This research is supported by the National Science and Technology Council, Taiwan, under grant No. MOST 110-2112-M-001-068-MY3, NSTC 113-2112-M-001-028-, 114-2112-M-001-012, 114-2811-M-001-094, and Academia Sinica, Taiwan, under a career development award under grant No. AS-CDA-111-M04. This research was supported in part by grant NSF PHY-2309135 to the Kavli Institute for Theoretical Physics (KITP) and grant NSF PHY-2210452 to the Aspen Center for Physics. KC acknowledges the support of the Alexander von Humboldt Foundation. Our computing resources were supported by the National Energy Research Scientific Computing Center (NERSC), a U.S. Department of Energy Office of Science User Facility operated under Contract No. DE-AC02-05CH11231 and the TIARA Cluster at the Academia Sinica Institute of Astronomy and Astrophysics (ASIAA).
Appendix A Classification of evolutionary pathways
@@
If $R_{P\rm max}$ and $R_{\rm max}$ differ by less than 1%, this indicates that the maximum radius is reached while the central pressure is still increasing, which implies that the core is still contracting at that time. For stars that satisfy this condition and have an effective temperature $\log(T_{\rm eff}/{\rm K})<3.8$ at $R_{\rm max}$, we classify them as RSG-C.